# Concept camera for the next-generation mm-wave cosmological surveys

### v1

@@ Abstract
Past millimeter-wave galaxy surveys have probed the brightest starburst galaxies only and suffered heavily from confusion. The interpretation of existing surveys has also been hindered by the lack of reliable redshift indicators for measuring distances for the entire sample. Thanks to recent advances in mm-wave detector technologies we can now overcome these limitations, and conduct the first truly volumetric surveys of star-forming galaxies at mm-wavelengths down to the L∗ luminosities of typical galaxies, with 1000 redshift slices spanning most of the Cosmic star-forming volume (1–12) with nearly uniform mass and luminosity selection. We describe an instrument concept capable of delivering such surveys with the technologies available today, which can be built and operated on a ground-based mm-wave facility in the near future. Such spectrometer cameras can resolve and redshift identify up to to 25,000 star-forming galaxies per year even when operated on a 10-m class telescope. On a larger aperture it can do the same faster or probe even deeper. We propose a loose, open-source collaboration to design, build, and operate one or several such cameras through the shared contributions of leading experts and telescopes from around the globe.
@@
far-infrared instrumentation, integral-field spectroscopy, cosmological star-formation survey, volumetric galaxy survey, line intensity mapping, cosmic star-formation history, imaging spectroscopy1 INTRODUCTION
@@
Stars form in dusty environments, obscured from direct view at the optical wavebands[1, 2]. The optical/UV light from the young stellar populations is absorbed by the dust, and then re-emitted as cold (35 K) thermal radiation in the far-infrared (FIR) bands, peaking near 100 m in the rest frame[3]. Even the Milky Way – with modest star-formation rates of 1 M⊙/year – radiates about half of its total bolometric luminosity in the FIR bands as a result, while galaxies with significantly higher star-formation rates emit over 99% of the luminosities at FIR wavelengths, and are difficult to spot optically, especially at higher redshifts.
@@
Cosmic star formation has peaked around redshift 2–3[4, 5]. The bulk of the stars we see today formed during intense starbursts in that era, when galaxies would have been at their most luminous (thanks to short-lived populations of massive OB stars), yet largely obscured at optical wavelengths. To catch galaxies during their intense episodes of star-formation, we really must study them at FIR wavelengths.
@@
Due to a strongly negative -correction on the Rayleigh-Jeans side of the thermal spectral energy distribution (SED), a star-forming galaxy will appear similarly bright at 1 mm wavelength regardless of its distance at 1–12 (Fig. 1). This flat mass-luminosity selection[3, 6] allows us to conduct unbiased luminosity-limited surveys of star-formation in the mm-band that spans much of the Cosmic history in which galaxies formed and evolved. This effect has been exploited for a number of FIR galaxy surveys[7, 8, 9]. The Herschel Multi-tiered Extragalactic Survey (HerMES)[8] alone has detected on the order of a 100,000 FIR galaxies with its long-wavelength (200–500 m) SPIRE[10] instrument.
@@
However, the flat selection at mm-wavelength is a mixed blessing. Because every line of sight can see star-forming galaxies at all distances equally, the detections alone do not inform us about the redshift of these objects; and the fields quickly get crowded (confused) as extremely bright galaxies fill up every beam on sky well before one could hope to individually detect the fainter, more typical Milky-Way-like galaxy populations. Attempts to use photometry as a way to guesstimate redhifts for these galaxies have proven ineffective at best[11], while spectroscopic follow-ups often rely on heavily biased cross-identifications[12] and can target at most a tiny fraction of the galaxies detected by the continuum surveys overall.
@@ 2 FUTURE SURVEYS
One way to address the confusion is to conduct surveys with larger telescope apertures. Increasing of aperture diameter by a factor allows detecting more sources before the confusion limit is reached. And with SMG number number counts typically for the (bright) submillimeter galaxy (SMG) populations[7], the confusion flux limit (i.e. mass-luminosity selection threshold) will scale inversely as . So larger apertures can help probe a little deeper, but they cannot magically solve the redshift identification issue.
@@
The CO rotational ladder offers perhaps the best tool for redshift identification. The lower- CO transitions are typically thermalized with a spectrum on the Rayleigh-Jeans side of their pseudo-blackbody spectral energy distributions (SEDs). SMGs are known to exhibit thermalized emission into the morerately high- (10) transitions[13, 14]. The observed spacing of the CO ladder is 115.3 GHz for a source at redshift . Thus, the number of transitions detectable in a broad-band spectrometer scales as at higher , while the measurement noise from co-adding fluxes from the necessary channels increases as . A such, the detectability of redshift, through the combination of multiple CO transitions in band, benefits from a negative -correction also with an effective emissivity index due to the multiplicity of available transitions.
The ultimate result of which is that CO redshifts have a similarly flat mass-luminosity selection thresholds at millimeter wavelengths as the dust continuum (Fig. 1). CO emission is also highly correlated with the thermal dust continuum in extragalactic star-forming environments[15], both being equally effective tracers of the star-forming gas. As such, there is no fundamental need to detect SMGs by their far-infrared continuum emission separately, when detection via the CO ladder alone yields effectively the same mass-luminosity measure for star-forming sources. (Arguably, the CO selects star-formation more purely, since the mm-wave continuum may be contaminated by synchrotron emission, such as from active galactic nuclei [AGNs].)
@@ 2.1 Volumetric Galaxy Surveys
Because the CO ladder from a distant galaxy leaves a unique spectral imprint as a function of redshift, an 200–1000 spectrometer covering octave-scale bandwidths can effectively sample 1000–5000 redshift slices above 230 GHz. Here is the minimum redshift at which at least 2 CO transitions are detected in the observable bandwidth. Even if two galaxies overlap spatially, they will leave distinct spectral imprints so long as they are separated by in the rest-frame velocity space (or in redshift space). That alone reduces confusion dramatically when compared to continuum surveys with comparable beam sizes. As a result, an 200–1000 survey instrument could detect up to 100–500 times more sources per beam than a continuum camera operated at the same wavelength before the spectral confusion limit is reached. Hence, an 200–1000 survey can probe star-forming galaxies to a mass / luminosity threshold an order-of-magnitude below that of a continuum survey at the same wavelengths for the same-sized telescope.
To put it into perspective, an 200 spectroscopic survey at 1.2 mm on a 10-m class telescope, would be confused at a mass / luminosity threshold that is 20 times below what Herschel/SPIRE could probe at its longest wavelength 500 m channel, while also peering further back in time. Consequently, it would resolve not only the most extreme ULIRG / HLIRG populations from before, but could for the first time systematically dip into the L∗ populations[16] that dominate the makeup of galaxies overall – while providing an unambiguous redshift identification for every single detected source also. On a 30-m aperture it could reach a factor of 2 deeper still.
@@ 2.2 Line Intensity Mapping (LIM)
A related application, for the same or very similar spectrometer camera, is Line Intensity Mapping (LIM)[17]. In this case, the goal is not the detection and unique redshift identification of individually resolved galaxies, but rather the statistical (unresolved) detection of entire populations, and the statistical (e.g. 3D) correlations that characterize them. Because LIM targets galaxies with similar properties (e.g. velocity dispersion) as the above described survey, its requirements are inherently similar, specifically with 200-1000 spectroscopic mapping[18]. However, because the goal is not the individual detection of sources, LIM surveys probe much larger areas albeit at shallower depths.
While deep galaxy surveys that can be performed effectively in a single atmospheric window (e.g. at 1.2 mm), LIM is best conducted over as much total bandwidth as possible, covering as many molecular and/or hyperfine transitions as possible. Thus, the ultimate LIM survey would require the next-level camera upgrade, featuring simultaneous multi-band capability. An ideal mm-wave LIM camera would cover the 3 mm, 2 mm, 1.2 mm, and 850 m atmospheric windows with fully populated focal plane spectrometers simultaneously. As such it may contain up to four separate focal-planes, with a correspondingly higher total detector count to match. A multi-band implementation will inevitably require a far more complicated optical and cryogenic configuration than what a single-band camera can get away with. Sufficient stray-light and out-of-band light rejection will be significantly more challenging to achieve. Nevertheless, it is possible with the right amount of ambition if one really wants it.
Both a single-band and multi-band spectroscopic camera could satisfy both science goals (not at the same time of course). It’s a question on which it is designed to do more optimally. The resolved galaxy surveys and LIM are inherently complementary to one another, both in their survey modes (small-deep vs. large-shallow), and their scientific outputs (individually resolved galaxies vs. statistically identified unresolved populations). Ideally, we will be pursuing both science goals in the not too distant future.
2.3 Enabling Technologies
Moderate resolution 200–1000 spectroscopic surveys are a compelling way forward. They promise to overcome both limitations of past (sub)millimeter continuum surveys, with significantly lesser confusion and direct, unambiguous, in-band redshift identifications. To conduct such surveys, however, we would need multi-beam spectroscopic cameras with moderate spectral resolutions of 200–1000 at mm-waves, preferably fully populating the field-of-view of a designated 10-m class or larger telescope. That in turn requires the following enabling technologies (scoped for a ground-based instrument, operated at 1 mm, with 3–10):
@@ 2.3 Enabling Technologies
moderate resolution spectral channelizing in a footprint that is comparable or smaller than the size of a focal-plane pixel (1 cm2). 111Alternatively, a set of slits and gratings illuminating a 2D array of detectors could in principle also be used to provide an 200 spectrometer with a 2D field of view also, expanding on the design of HIRMES[19] but with an additional tiling of slits and gratings, and hence spectra. However, we consider that approach both more complex and less proven, and hence we do not consider it here in detail.
@@
direct detectors with NEP310-18 W/Hz1/2, with footprint 1 mm2 to allow packing 200 detectors onto a focal-plane pixel.
@@
Recent advances in technology have readily provided us with all three prerequisites. On-chip mm-wave spectrometers, like SuperSpec[20, 21] or DESHIMA[22, 23], with 100–1000 have been demonstrated. Crucially, the channelizer part of these is already much smaller than the typical 1 cm2 of a mm-wave focal-plane pixel of an envisaged mm-wave camera (most of the real-estate taken up by the readout resonators of the KIDs for both). Highly sensitive KIDs that far surpass the NEP requirement for background-limited operation have been produced[24], and KIDs that meet the size and multiplexing requirements are presently being deployed[25, 26]. All that remains is combining these advances in the form of a new mm-wave integral field spectrometer (IPS) camera that can fill the field-of-view (FoV) of a 10-m class or larger mm-wave telescope. It could be done ’today’.
@@ 3 INSTRUMENT CONCEPT
Spectroscopic galaxy surveys at mm-wavelengths could be conducted from ground-based facilities. The 1.2 mm atmospheric window is especially promising in terms of yielding the highest detection rates, through a combination of factors: (1) the widest fractional bandwidth observable from the ground (0.5), which also results in the (2) lowest cutoff, (3) low atmospheric opacities most of the time even at moderate altitudes, and (4) allows for cm2-scale packed pixel footprints, and (5) a reasonably flat CO redshift selection for 1.5–12, being firmly on the Rayleigh-Jeans side of the SED even for sources of extreme redshift (6). However, we’ll consider designs for the neighboring 2 mm and 850 m atmospheric windows also. See Tables 1–3 for details.
@@
2 | 125 – 175 | 50 | 0.31 | 0.5 | 200 |
1.2 | 190 – 310 | 120 | 0.15 | 0.7 | 280 |
0.87 | 330 – 360 | 30 | 2.9 | 0.13 | 52 |
We baseline the camera concept for a 10-m class telescope. These are the smallest of the ’large’ mm-wave telescopes, which abound thanks to many ALMA prototypes (e.g. APEX[27], GLT[28]), as well as existing telescopes of other designs (e.g., LCT[29] or SPT[30]). The point is to show that a 10-m class telescope is perfectly capable delivering the transformative science when equipped with a background-limited 200 spectrometer camera, and operated from a site where PVW2 mm is typical. Larger telescopes (e.g. JCMT, IRAM 30-m[31], LMT[32], or AtLAST[33]) could do the same thing faster and/or go deeper still. A 50-m class telescope may also provide astrometry of the sources at the 1–2′′ level, which is sufficient for direct follow-up in most other bands.
We will assume a 10 arcmin FoV, which is achievable for most 10-m class mm-wave telescopes (e.g. in the Cassegrain cabin of APEX prototypes) without major modifications. Telescopes that offer a smaller field-of-view are not ruled out either – but their mapping speeds would be reduced proportionally to the area that can be imaged instantaneously. (This penalty, however, may be more than compensated by gains from a larger aperture size, e.g. for a 30-m telescope). Conversely, a larger field-of-view can increase the speed and yield of surveys, provided that it can be populated with the necessary number of detectors also.
@@
2 | 86 | 0.55 | 6.8 |
1.2 | 220 | 1.1 | 11 |
0.87 | 450 | 1.9 | 16 |
The in-pixel 200 channelizing of the incident radiation necessitates an antenna-coupled design, whereby light is coupled into a lithographic transmission line filterbank. Because of the relatively narrow-band (200) flux measurement, we want to focus as much of the incident light from a point-source onto a single pixel to achieve the best sensitivities. The incident radiation may be focused on the antenna by a suitable feedhorn or a lens. The most efficient coupling to the telescope beam is likely achieved with feedhorns packed at a 2 spacing. To maximize sensitivity, ideally we couple both orthogonal linear polarizations for a given channel into the same total-power detector. Such dual-polarization driving ought to be possible for kinetic inductance detectors (KIDs), where a single detector could absorb power from two separate channelizer structures originating from orthogonally polarized antenna inputs. The dual-polarization coupling would be a novelty, but it is also an non-essential. Without it, the effective survey speed will be reduced by a factor of 2, but the survey data quality will be unaffected otherwise, since the CO emission is not expected to be strongly polarized.
Given the above parameters (10-m telescope, 10 arcmin FoV, 1.2 mm, =5, 2 pixel spacing), a suitable spectrometer camera would have around 100 pixels, arranged in a hexagonal pattern at spacing, fully populating the available field-of-view. Each pixel, with a physical area of 1.1 cm2, must fit an 200 channelizer for 280 Nyquist sampled channels, and a matching number of detectors per pixel, or 25 thousand detectors in all. Like for SuperSpec[21], the channelizer can be provided via thin-film microstrip transmission lines, on an area that is 20 m wide and long. For the 1.2 mm band that translates to an area of 0.5 mm2 assuming a suitable dielectric with 10. Accordingly, around 99% of the focal-plane real-estate can be used by the 280 detectors per pixel, which sets the size of individual detectors to 0.44 mm2 (or equivalently [0.66 mm]2). That detector area is not too different from that of the KIDs currently being deployed in the high-frequency array (HFA) of the APEX MKID camera[25, 26].
@@
2 | 42 | 37 | 7,400 | 1.52 | 0.82 | 10,800 |
1.2 | 25 | 91 | 25,480 | 0.44 | 0.45 | 23,300 |
0.87 | 18 | 169 | 8,788 | 1.29 | 0.30 | 5,500 |

### v2

# Concept integral field unit spectrometer instrument for the next-generation mm-wave cosmological surveys

@@ Abstract
Past millimeter-wave galaxy surveys have primarily probed the brightest starburst galaxies only and suffered heavily from confusion. The interpretation of existing surveys has also been hindered by the lack of reliable redshift indicators for measuring distances for entire samples. Thanks to recent advances in mm-wave detector technologies we can now overcome these limitations, and conduct the first truly volumetric surveys of star-forming galaxies at mm-wavelengths approaching the luminosities of typical galaxies, with 1000 redshift slices spanning most of the Cosmic star-forming volume (1–12) with nearly uniform mass and luminosity selection. We describe an instrument concept capable of delivering such surveys with the technologies available today, which can be built and operated on a ground-based mm-wave facility in the near future. Such integral field unit spectrometers can resolve and identify redshifts for up to to 25,000 star-forming galaxies per year even when operated on a 10-m class telescope. On a larger aperture it can do the same faster or probe even deeper. We propose a collaboration open-source initiative to design, build, and operate one or several such cameras through the shared contributions of leading experts and telescopes from around the globe.
@@
[…]
1 Introduction
@@
Stars form in dusty environments, obscured from direct view at the optical wavebands[Devlin2009, Puget2006]. The optical/UV light from the young stellar populations is absorbed by the dust, and then re-emitted as cold (35 K) thermal radiation in the far-infrared (FIR) bands, peaking near 100 m in the rest frame[Kovacs2010]. Even the Milky Way – with modest star-formation rates of 1 M⊙/year – radiates about half of its total bolometric luminosity in the FIR bands as a result, while galaxies with significantly higher star-formation rates, such as Arp220, emit over 99% of the luminosities at FIR wavelengths, and are difficult to spot optically, especially at higher redshifts.
@@
Cosmic star formation has peaked around redshift 2–3[Chapman2003, Madau2014]. The bulk of the stars we see today formed during intense starbursts in that era, when galaxies would have been at their most luminous phase (thanks to short-lived populations of massive OB stars), yet largely obscured at optical wavelengths. Catching galaxies during their intense episodes of star-formation requires FIR/(sub)mm surveys.
@@
Due to a strongly negative -correction on the Rayleigh-Jeans side of the thermal spectral energy distribution (SED), a star-forming galaxy will appear similarly bright at 1 mm wavelength regardless of its distance at 1–12 (Fig. 1). This flat mass-luminosity selection[Kovacs2010, Staguhn2014] allows us to conduct unbiased luminosity-limited surveys of star-formation in the mm-band that spans much of the Cosmic history in which galaxies formed and evolved. This effect has been exploited for a number of FIR galaxy surveys[Weiss2009, Oliver2012, Simpson2019]. The Herschel Multi-tiered Extragalactic Survey (HerMES)[Oliver2012] alone has detected on the order of a 100,000 FIR galaxies with its long-wavelength (200–500 m) SPIRE[SPIRE] instrument.
@@
However, the flat selection at mm-wavelength is a mixed blessing. Because every line of sight can see star-forming galaxies at all distances equally, the detections alone do not inform us about the redshift of these objects; and the fields quickly get crowded (confused) as extremely bright galaxies fill up every beam on sky well before one could hope to individually detect the fainter, more typical Milky-Way-like galaxy populations. Attempts to use photometry as a way to estimate redshifts for these galaxies have proven ineffective at best[Cox2023], while optical/NIR spectroscopic follow-up often relies on heavily biased cross-identifications[Pope2006] and can target at most a tiny fraction of the galaxies detected by the continuum surveys overall.
In this paper, we demonstrate that a 100-pixel integral field unit (IFU) spectrometer, with 200 resolution, can fill the focal planes of many existing (or future) 10-m class (or larger) mm-wave telescopes, and cover the entire 1 mm atmospheric window with 20,000 detectors. The technologies required for such an instrument exist ’today’ and could enable volumetric galaxy surveys, down to around L∗ luminosities efficiently, yielding up to 23,000 galaxies per survey year. In Section 2 we describe future volumetric surveys including the possibility of Line Intensity Mapping (LIM). In Section 3 we list some of the technologies that enable an IFU spectometer ’today’. In Section 4 we describe what a typical realization for an existing 10-m class telescope may entail, followed by Section 5 on how the designs may be adjusted / scoped for other telescopes, technologies, or resources. Finally, to build such an instrument for the Greenland Telescope (GLT), and to facilitate parallel development of similar instrument we suggest and open-source collaboration in Section 6.
@@ 2 Future Surveys
One way to address the confusion is to conduct surveys with larger telescope apertures. Increasing of aperture diameter by a factor allows detecting more sources before the confusion limit is reached. And with SMG number counts typically for the (bright) submillimeter galaxy (SMG) populations[Weiss2009], the confusion flux limit (i.e. mass-luminosity selection threshold) will scale inversely as . So larger apertures can help probe a little deeper, but they cannot effectively solve the redshift identification issue.
@@
The CO rotational ladder offers perhaps the best tool for redshift identification. The lower- CO transitions are typically thermalized with a spectrum on the Rayleigh-Jeans side of their pseudo-blackbody spectral energy distributions (SEDs). SMGs are known to exhibit quasi-thermalized emission into the moderately high- (10) transitions[Weiss2007, Boogaard2020]. The observed spacing of the CO ladder is 115.3 GHz for a source at redshift . Thus, the number of transitions detectable in a broad-band spectrometer scales as at higher , while the measurement noise from co-adding fluxes from the necessary channels increases as . As such, the detectability of redshift, through the combination of multiple CO transitions in band, benefits from a negative -correction also with an effective emissivity index due to the multiplicity of available transitions.
The ultimate result of which is that CO redshifts have a similarly flat mass-luminosity selection thresholds at millimeter wavelengths as the dust continuum (Fig. 1). CO emission is also highly correlated with the thermal dust continuum in extragalactic star-forming environments[Weiss2008], both being equally effective tracers of the star-forming gas. As such, there is no fundamental need to detect SMGs by their far-infrared continuum emission separately, when detection via the CO ladder alone yields effectively the same mass-luminosity measure for star-forming sources. (Arguably, the CO selects star-formation more purely, since the mm-wave continuum may be contaminated by synchrotron emission, such as from active galactic nuclei [AGNs].)
@@ 2.1 Volumetric Galaxy Surveys
Because the CO ladder from a distant galaxy leaves a unique spectral imprint as a function of redshift, an 200–1000 spectrometer covering octave-scale bandwidths can effectively sample 1000–5000 redshift slices above 230 GHz. Here is the minimum redshift at which at least two CO transitions are detected in the observable bandwidth. Even if two galaxies overlap spatially, they will leave distinct spectral imprints so long as they are separated by in the rest-frame velocity space (or in redshift space). That alone reduces confusion dramatically when compared to continuum surveys with comparable beam sizes. As a result, an 200–1000 integral field unit spectrometer instrument could detect up to 100–500 times more sources per beam than a continuum camera operated at the same wavelength before the spectral confusion limit is reached. Hence, an 200–1000 survey can probe star-forming galaxies to a mass / luminosity threshold an order-of-magnitude below that of a continuum survey at the same wavelengths for the same-sized telescope.
To put it into perspective, an 200 spectroscopic survey at 1.2 mm on a 10-m class telescope[GLT-survey], would be confused at a mass / luminosity threshold that is 20 times below what Herschel/SPIRE could probe at its longest wavelength 500 m channel, while also peering further back in time. Consequently, it would resolve not only the most extreme ULIRG / HLIRG populations from before, but could for the first time systematically dip into the populations[Cooray2005] (cf. Fig. 1 right panel) that dominate the makeup of galaxies overall – while providing unambiguous redshift identification for every single detected source also. On a 30-m aperture it could reach about twice as deep still. Such surveys have been proposed for future 50-m class telescopes[Kohno2018, Geach2019], but not for the existing 10-m class telescopes, which could deliver significant results sooner.
@@ 2.2 Line Intensity Mapping (LIM)
A related application, for the same or very similar IFU spectrometer, is Line Intensity Mapping (LIM)[Azadeh2022, GLT-LIM, Vallini2023]. In this case, the goal is not the detection and unique redshift identification of individually resolved galaxies, but rather the statistical (unresolved) detection of entire populations, and the statistical (e.g. 3D) correlations that characterize them. Because LIM targets galaxies with similar properties (e.g. velocity dispersion) as the above described survey, its requirements are inherently similar, specifically with 200-1000 spectroscopic mapping[Karkare2018, Karkare2022]. However, because the goal is not the individual detection of sources, LIM surveys probe much larger areas albeit at shallower depths.
While deep galaxy surveys that can be performed effectively in a single atmospheric window (e.g. at 1.2 mm), LIM is best conducted over as much total bandwidth as possible, covering as many molecular and/or hyperfine transitions as possible. Thus, the ultimate LIM survey would require the next-level instrument upgrade, featuring simultaneous multi-band capability. An ideal mm-wave LIM instrument would cover the 3 mm, 2 mm, 1.2 mm, and 850 m atmospheric windows with fully populated focal plane spectrometers simultaneously. As such it may contain up to four separate focal-planes, with a correspondingly higher total detector count to match. A multi-band implementation will inevitably require a far more complicated optical and cryogenic configuration than what a single-band IFU spectrometer can get away with. Sufficient stray-light and out-of-band light rejection will be significantly more challenging to achieve, and it requires more detectors to be read out. Nevertheless, a multiband IFU spectrometer is also possible with the right amount of ambition.
Both a single-band and multi-band spectroscopic instrument could satisfy both science goals (e.g. time-sharing to conduct two distinct surveys side-by-side). It is a matter of which type of science the instrument is designed to do more optimally. The resolved galaxy surveys and LIM are inherently complementary to one another, both in their survey modes (small-deep vs. large-shallow), and their scientific outputs (individually resolved galaxies vs. statistically identified unresolved populations). Ideally, we will be pursuing both science goals in the not too distant future.
3 Enabling Technologies
Moderate resolution 200–1000 spectroscopic surveys are a compelling way forward. They promise to overcome both limitations of past (sub)millimeter continuum surveys, with significantly lesser confusion and direct, unambiguous, in-band redshift identifications. To conduct such surveys, however, we would need multi-beam spectrometers with moderate spectral resolutions of 200–1000 at mm-waves, preferably fully populating the field-of-view of a designated 10-m class or larger telescope. That in turn requires the following enabling technologies (scoped for a ground-based instrument, operated at 1 mm, with 3–10):
@@ 3 Enabling Technologies
Moderate resolution spectral channelizing in a footprint that is comparable or smaller than the size of a focal-plane pixel (1 cm2). By integrating an entire spectrometer into the footprint of a focal-plane pixel, it becomes possible to create arbitrarily large, densely-packed, and scalable IFU spectrometer arrays.
@@
Direct detectors with NEP310-18 W/Hz1/2, with footprint 1 mm2 to allow packing 200 (or more) detectors onto a focal-plane pixel.
@@
Recent advances in technology have readily provided us with all three prerequisites, separately. On-chip mm-wave spectrometers, like SuperSpec[SuperSpec-concept, SuperSpec], DESHIMA[DESHIMA, DESHIMA2], or SPT-SLIM[SPT-SLIM], with of a few hundred have been demonstrated (albeit in considerably larger footprints than what we aim for). Crucially, the channelizer part of these is already much smaller than the typical 1 cm2 of a mm-wave focal-plane pixel of an envisaged mm-wave instrument (most of the real-estate taken up by the readout resonators of the KIDs for both). Highly sensitive KIDs that far surpass the NEP requirement for background-limited operation have been produced[SPACEKIDS], and KIDs that meet the size and multiplexing requirements are presently being deployed[AMKID, AMKID-new] (albeit not as antenna-coupled spectrometer devices). What remains is combining these advances in the form of a new mm-wave integral field unit spectrometer that can fill the field-of-view (FoV) of a 10-m class or larger mm-wave telescope with densely packed spectrometer pixels. It could be done ’today’. This view is reinforced by the current development of TIFUUN[TIFUUN] for ASTE[ASTE] to realize a very similar IFU spectrometer to the one we propose here.
@@ 4 Instrument Concept
Spectroscopic galaxy surveys at mm-wavelengths could be conducted from ground-based facilities. Our posit is that a 100-pixel IFU spectrometer with 200 could fill a 10′ field-of-view of a 10-m class mm-wave facility, covering the entire 1.2 mm atmopsheric window, and would detect up to 20,000 redshift complete star-forming galaxies down to near luminosities in 1000 h of operation.
The 1.2 mm atmospheric window is especially promising in terms of yielding the highest detection rates, through a combination of factors: (1) the widest fractional bandwidth observable from the ground (0.5), which also results in the (2) lowest cutoff, (3) low atmospheric opacities most of the time even at moderate altitudes, and (4) allows for cm2-scale packed pixel footprints, and (5) a reasonably flat CO redshift selection for 1.5–12, being firmly on the Rayleigh-Jeans side of the SED even for sources of extreme redshift (6). However, we’ll consider designs for the neighboring 2 mm and 850 m atmospheric windows also. See Tables 1–3 for details.
@@
| 2 | 125 – 175 | 50 | 0.31 | 0.5 | 200 |
| 1.2 | 190 – 310 | 120 | 0.15 | 0.7 | 280 |
| 0.87 | 330 – 360 | 30 | 2.9 | 0.13 | 52 |
We baseline the IFU spectrometer concept for a 10-m class telescope. These are the smallest of the ’large’ mm-wave telescopes, which abound thanks to many 12-m ALMA prototypes (e.g. APEX[APEX], GLT[GLT]), as well as existing telescopes of other designs (e.g., LCT[LCT] or SPT[SPT], ASTE[ASTE]). The point is to show that a 10-m class telescope is perfectly capable delivering the transformative science when equipped with a background-limited 200 IFU spectrometer, and operated from a site where PVW2 mm is typical.
We will assume a 10 arcmin FoV, which is achievable for most 10-m class mm-wave telescopes (e.g. in the Cassegrain cabin of the ALMA prototypes) without major modifications.
@@
| 2 | 86 | 0.55 | 6.8 |
| 1.2 | 220 | 1.1 | 11 |
| 0.87 | 450 | 1.9 | 16 |
The in-pixel 200 channelizing of the incident radiation necessitates an antenna-coupled design, whereby light is coupled into a lithographic transmission line filterbank. Because of the relatively narrow-band (200) flux measurement, we want to focus as much of the incident light from a point-source onto a single pixel to achieve the best sensitivities. The incident radiation may be focused on the antenna by a suitable feedhorn or a lens. The most efficient coupling to the telescope beam is likely achieved with feedhorns packed at a 2 spacing. To maximize sensitivity, ideally we couple both orthogonal linear polarizations for a given channel into the same total-power detector. Such dual-polarization driving ought to be possible for kinetic inductance detectors (KIDs), where a single detector could absorb power from two separate channelizer structures originating from orthogonally polarized antenna inputs.
Given the above parameters (10-m telescope, 10 arcmin FoV, 1.2 mm, =5, 2 pixel spacing), a suitable IFU spectrometer would have around 100 pixels, arranged in a hexagonal pattern at spacing, fully populating the available field-of-view. Each pixel, with a physical area of 1.1 cm2, must fit an 200 channelizer for 280 Nyquist sampled channels, and a matching number of detectors per pixel, or 25 thousand detectors in all. Like for SuperSpec[SuperSpec], the channelizer can be provided via thin-film microstrip transmission lines, on an area that is 20 m wide and long. For the 1.2 mm band that translates to an area of 0.5 mm2 assuming a suitable dielectric with 10. Accordingly, around 99% of the focal-plane real-estate can be used by the 280 detectors per pixel, which sets the size of individual detectors to 0.44 mm2 (or equivalently [0.66 mm]2). That detector area is not too different from that of the KIDs currently being deployed in the high-frequency array (HFA) of the APEX MKID camera[AMKID, AMKID-new].
@@
| 2 | 42 | 37 | 7,400 | 1.52 | 0.82 | 10,800 |
| 1.2 | 25 | 91 | 25,480 | 0.44 | 0.45 | 23,300 |
| 0.87 | 18 | 169 | 8,788 | 1.29 | 0.30 | 5,500 |
Table 3 summarizes the expected performance and scientific yield of a IFU spectrometer with 10 arcmin FoV on a 10-m telescope for 1 year (1000 hours on source) of operation for the deep galaxy survey. Table 4 does the same for the line intensity mapping use case, assuming a 10 deg2 survey area. For the galaxy survey we show the confusion-limited survey area (i.e., mapping speed) and the expected number () of individually detected, redshift-complete galaxies, per year of operation. For LIM we identify the line species, which are expected to dominate the power spectrum signal, while noting also that a multitude of lines from different redshifts will also contribute to the autocorrelation power measured in each band. Shown are the estimated noise per mode (), the total Number of modes measured (), and the lowest wavenumber nominally measured by the survey (), assuming no limitations from atmospheric fluctuations or other sources of 1/ noise.
| Line | |||||
| (mm) | (K2 Mpc3 / ) | ( 106) | ( / Mpc) | ||
| 2 | CO(3-2) | 1.3054 | 13,000 | 5.4 | 0.0399 |
| 1.2 | CO(4-3) | 0.8443 | 24,000 | 2.7 | 0.0485 |
| 0.87 | CO(6-5) | 1.0047 | 58,000 | 7.3 | 0.0482 |